A particle moves along the side AB, BC, and CD, of a square of side 25m with a velocity of 15ms-1. Its average velocity is

(a) 15ms-1

(b) 10s-1

(c) 7.5ms-1

(d) 5ms-1

Grade:9

1 Answers

vikas askiitian expert
13 years ago

time taken by particle to cross any side = t

speed = d/t               (d is length of side)

t = 25/15 ..............1

now when particle moves from AB,BC,CD then it reaches to adjecent vertex D of square ...

total time taken = T = 3t =3(25/15) = 5sec

now average velocity = total displacement/total time

displacement BW initial poiny A & D is 25m so

average velocity  =  25/5 = 5m/s

option d is correct